A ball (i.e., a hollow sphere), sliding on a horizontal friction less surface at 20 m/s, encounters a rough surface and begins to roll. Assuming that energy is conserved what is the speed of the ball after it begins to roll?
Answer #1

0.5 m v^2 = 0.5 mu^2 + 0.5 I w^2

m v^2 = m u^2 + (2/3) m r^2 u^2/r^2

20^2 = u^2 + (2/3) u^2

400 = 1.67 u^2

u = 15.49 m/s

Comment in case any doubt please rate my answer....